OpenStudy (experimentx):

Show that \[ \sum_{n=1}^\infty {\sin n \over n} = {-1 + \pi \over 2}\]

OpenStudy (anonymous):

it has to do something with fourier series...or maybe there is an easier way

OpenStudy (anonymous):

\[ \sum_{n=1}^\infty {\sin n \over n} = {-1 + \pi \over 2}\]

OpenStudy (experimentx):

yep!!

OpenStudy (experimentx):

can't help myself with any other method ... few days ago i tried fooling around ... i couldn't do it.

OpenStudy (anonymous):

this guy did a good work here http://answers.yahoo.com/question/index?qid=20101031215404AAQmZGn

OpenStudy (experimentx):

there one here http://math.stackexchange.com/questions/13490/proving-that-the-sequence-f-nx-sum-limits-k-1n-frac-sinkxk-is don't understand this nasty third step

OpenStudy (experimentx):

looks this yahoo is worth try !! i haven't seen that one.

OpenStudy (anonymous):

but there is hole in yahoo proof .... is this converges? \[\sum_{n=1}^{\infty} e^{inz}\] for \(|z|<1\) and \(z=1\)

OpenStudy (anonymous):

if converges for \(|z|<1\) we must go to Abel's theorem for \(z=1\)

OpenStudy (experimentx):

wonskian =/= 0 ??

OpenStudy (experimentx):

this is  this might converge (let's say it for now) the other one e^(inx) diverges (or not known) ... i guess

OpenStudy (experimentx):

hold on let's do this